Given: x = csct and y = cott. Find y’.


1
Expert's answer
2022-02-24T11:23:34-0500

"y'=\frac{dy}{dx}=\frac{\frac{dy}{dt}}{\frac{dx}{dt}}=\frac{}{}=\frac{-\csc^2t}{-\cot t\csc t}=\sec t."